OpenStudy (anonymous):

Evaluate the function at each specified value of the independent variable and simplify. f(x-6)=|(X-6)|/(x-6)

OpenStudy (anonymous):

f(x-6)=|(X-6)|/(x-6) |x-6| will be = x-6 if x-6>=0 => x>=6 |x-6| will be -(x-6) if x-6<0 => x<6 hence f(x-6)=(x-6)/(x-6) when x>=6 and f(x-6)=-(x-6)/x-6 when x<6 hence f(x-6)=1 when x>=6 and f(x-6)=-1 when x<6